# Assimilative Causal Inference

### v1

@@ A.2.3 Filter and Smoother
Under mild spatial regularity assumptions on the parameters in (1.1), both the filter and smoother distributions are absolutely continuous with respect to the Lebesgue measure on [21]. This guarantees the existence of their probability density functions (PDFs) for all $t\in[0,T]$, which we define as:
Filter: $p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x}):=p\big{(}\mathbf{y}(t)|\mathbf{x}(s\leq t)\big{)},$ (1.2)
Smoother: $p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}):=p\big{(}\mathbf{y}(t)|\mathbf{x}(s\leq T)\big{)}.$
@@ A.3 Explicit Expression of the ACI Framework for Gaussian Distributions
The relative entropy (also known as the Kullback-Leibler divergence) between two given PDFs $p(\mathbf{u})$ and $q(\mathbf{u})$ is defined as [13]:
@@
The relative entropy (1.3) benefits from a simple and explicit formula when the distributions $p$ and $q$ are real-valued Gaussian densities. Specifically, for $p(\mathbf{u})$ being the PDF of $\mathcal{N}_{M}(\boldsymbol{\mu}_{\text{p}},\mathbf{R}_{\text{p}})$ and $q(\mathbf{u})$ of $\mathcal{N}_{M}(\boldsymbol{\mu}_{\text{q}},\mathbf{R}_{\text{q}})$, then [3, 12]:
@@
A.4 Conditional ACI
@@
$\mathcal{P}\big{(}p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}_{\text{{A}}},\mathbf{x}_{\text{{B}}}),p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x}_{\text{{A}}},\mathbf{x}_{\text{{B}}})\big{)},\quad t\in[0,T].$ (1.7)
@@ A.4.1 Analyzing how $\mathbf{x}_{\text{{B}}}$ Affects the State Estimation of $\mathbf{y}(t)$
To analytically characterize how non-target variables $\mathbf{x}_{\text{B}}$ influence the $\mathbf{y}(t)$ estimation during Bayesian data assimilation, it suffices to examine the governing evolution equations for the unnormalized filter and smoother densities associated with (1.1) [21, 15]. Under sufficient regularity conditions, we have that the unnormalized filter density $u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})$ satisfies a linear random partial differential equation (PDE) for $0\leq t\leq T$:
$\mathrm{d}u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})=\underbrace{\mathcal{L}^{*}u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})\mathrm{d}t}_{\displaystyle\mathclap{\text{Forecast}}}+\underbrace{\mathcal{M}^{*}u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})\cdot\mathrm{d}\mathbf{x}}_{\displaystyle\mathclap{\text{Analysis}}},$ (1.8)
@@
$\mathcal{M}^{*}u_{t}(\mathbf{y}):=(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}\big{(}-\nabla_{\mathbf{y}}\cdot\big{(}(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}u_{t}\big{)}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}\mathbf{f}^{\mathbf{x}}u_{t}\big{)},$ (1.10)
@@
$\overleftarrow{\mathrm{d}u_{t}^{\text{{s}}}}(\mathbf{y}|\mathbf{x})=\underbrace{\mathcal{L}u_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})\mathrm{d}t}_{\displaystyle\mathclap{\text{Forecast}}}+\underbrace{\mathcal{M}u_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})\cdot\overleftarrow{\mathrm{d}\mathbf{x}}}_{\displaystyle\mathclap{\text{Analysis}}},$ (1.11)
@@
The diffusion interactions $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})$ are defined in the same manner as $\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}$, while $\overleftarrow{\mathrm{d}u_{t}^{\text{{s}}}}$ and $\overleftarrow{\mathrm{d}\mathbf{x}}$ denote backward stochastic Itô integrals (i.e., the negative of the usual differentials).
@@ A.4.3 Arguments in Favor of the Conditional ACI Framework
The generalized ACI framework presented in Section AA.4 and defined by (1.13)–(1.14) offers several key advantages for establishing (1.15):
@@
    Complete elimination of $\mathbf{x}_{\text{B}}$’s influence during data assimilation, ensuring any uncertainty reduction in $\mathbf{y}(t)$ stems solely from $\mathbf{x}_{\text{A}}$, consistent with our ACI objectives.
@@ A.4.4 Further Justification: Why Infinite Uncertainty Outperforms Marginalization
One might attempt to infer a conditional causal relationship from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$ by marginalizing over $\mathbf{x}_{\text{B}}$:
$p_{t}\big{(}\mathbf{y}\big{|}\mathbf{x}_{\text{A}}(s\leq T)\big{)}=\int_{\mathrm{supp}(\mathbf{x}_{\text{B}}(t))}p_{t}\big{(}\mathbf{y},\mathbf{x}_{\text{B}}\big{|}\mathbf{x}_{\text{A}}(s\leq T)\big{)}{\mathrm{d}}\mathbf{x}_{\text{B}}(t).$
@@
$\mathbf{x}_{\text{{A}}}$$\mathbf{x}_{\text{{B}}}$$\mathbf{y}$
For this causal chain, this marginalization (or conditioning) method might falsely suggest a direct causal link from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$. While these filter- and smoother-based distributions of $\mathbf{y}(t)$ given $\mathbf{x}_{\text{A}}$ may differ after marginalization (or conditioning), we cannot attribute causation to $\mathbf{x}_{\text{A}}$ under (1.4). The uncertainty reduction in $\mathbf{y}(t)$ inherently incorporates information transfer from $\mathbf{x}_{\text{B}}(t)$ due to their coupled dynamics, unlike the distributions in (1.13) where $\mathbf{x}_{\text{B}}$ trajectories are known but carry infinite likelihood uncertainty.
A.5 Instantaneous Discovery of the Causal Influence Range (CIR)
The relative entropy between the smoother and filter solutions in (1.4) quantifies both the existence and strength of an assimilative causal link from $\mathbf{y}(t)$ to $\mathbf{x}$. Similarly, the generalized metric in (1.14) measures the conditional causal link $\big{(}\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big{)}\big{|}\mathbf{x}_{\text{B}}$ and its intensity. However, these ACI metrics alone cannot determine the temporal extent of causal influence, namely, how many future values of $\mathbf{x}$ (or $\mathbf{x}_{\text{A}}$ conditioned on $\mathbf{x}_{\text{B}}$) are affected by $\mathbf{y}(t)$ over $[t,T]$. We define this temporal characteristic as the causal influence range (CIR), representing the future time window where $\mathbf{y}(t)$’s causal impact persists. Below we develop the theory for unconditional CIR; the conditional case (CCIR) follows analogously by substituting the appropriate equations with those from Section AA.4.
In chaotic turbulent systems, new observations of $\mathbf{x}$ influence $\mathbf{y}(t)$’s estimation only within finite time windows. More chaotic dynamics accelerate memory decay, typically exhibiting exponential decay modulated by factors like $\mathbf{x}$’s signal-to-noise ratio [2]. Our framework estimates this CIR duration for $\mathbf{y}(t)\rightarrow\mathbf{x}$ by determining how many future $\mathbf{x}$ values are meaningfully affected.

### v2

@@ A.2.3 Filter and Smoother
Under mild spatial regularity assumptions on the parameters in (1.1), both the filter and smoother distributions are absolutely continuous with respect to the Lebesgue measure on [rozovsky2012stochastic]. This guarantees the existence of their probability density functions (PDFs) for all $t\in[0,T]$, which we define as:
Filter: $p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})=p\big(\mathbf{y}(t)|\mathbf{x}(s\leq t)\big),$ (1.2)
Smoother: $p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x})=p\big(\mathbf{y}(t)|\mathbf{x}(s\leq T)\big).$
@@ A.3 Explicit Expression of the ACI Framework for Gaussian Distributions
The relative entropy (also known as the Kullback-Leibler divergence) between two given PDFs $p(\mathbf{u})$ and $q(\mathbf{u})$ is defined as [kleeman2011information]:
@@
The relative entropy (1.3) benefits from a simple and explicit formula when the distributions $p$ and $q$ are real-valued Gaussian densities. Specifically, for $p(\mathbf{u})$ being the PDF of $\mathcal{N}_{M}(\boldsymbol{\mu}_{\text{p}},\mathbf{R}_{\text{p}})$ and $q(\mathbf{u})$ of $\mathcal{N}_{M}(\boldsymbol{\mu}_{\text{q}},\mathbf{R}_{\text{q}})$, then [cai2002mathematical, kleeman2002measuring]:
@@
In Algorithm 1 we provide a brief outline of how to implement ACI in pseudocode.
Data: Observations of $\mathbf{x}$ in $[0,T]$ (continuous or discrete), $p_{0}^{\text{f}}(\mathbf{y}|\mathbf{x})$
Result: $\mathcal{P}\big(p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}),p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})\big)$ for $t\in[0,T]$, used to assess $\mathbf{y}(t)\rightarrow\mathbf{x}$ at each time
1
2for $0\leq t\leq T$ do
    Use (1.1) and the observations of $\mathbf{x}$ to compute $p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})$;
    /*/ E.g., see (1.8)--(1.10) for (1.1) or Theorem B.1 for the system in (2.1)
3
4 end for
5
6for $T\geq t\geq 0$ (moving backward) do
    Use $p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})$ and the observations of $\mathbf{x}$ to compute $p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x})$;
    /*/ E.g., see (1.11)--(1.12) for (1.1) or Theorem B.2 for the system in (2.1)
7
8 end for
9
10for $0\leq t\leq T$ do
    Calculate $\mathcal{P}\big(p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}),p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})\big)=\int p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x})\log\left(\displaystyle\frac{p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x})}{p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})}\right){\mathrm{d}}\mathbf{y}(t)$;
    /*/ Can also use any other type of information metric, e.g., a different $f$-divergence
11    if $\mathcal{P}\big(p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}),p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x})\big)>0$ then
12        $\mathbf{y}(t)\rightarrow\mathbf{x}$;
13
14   else
15       $\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}$;
16    end if
17
18 end for
Algorithm 1 ACI for the identification of $\mathbf{y}(t)\rightarrow\mathbf{x}$ over $t\in[0,T]$
A.4 Conditional ACI in the Presence of Additional Non-Target Variables
@@
$\mathcal{P}\big(p_{t}^{\text{s}}(\mathbf{y}|\mathbf{x}_{\text{{A}}},\mathbf{x}_{\text{{B}}}),p_{t}^{\text{f}}(\mathbf{y}|\mathbf{x}_{\text{{A}}},\mathbf{x}_{\text{{B}}})\big),\qquad t\in[0,T].$ (1.7)
@@ A.4.1 Analyzing how $\mathbf{x}_{\text{{B}}}$ Affects the State Estimation of $\mathbf{y}(t)$
To analytically characterize how non-target variables $\mathbf{x}_{\text{B}}$ influence the $\mathbf{y}(t)$ estimation during Bayesian data assimilation, it suffices to examine the governing evolution equations for the unnormalized filter and smoother densities associated with (1.1) [rozovsky2012stochastic, liptser2001statistics]. Under sufficient regularity conditions, we have that the unnormalized filter density $u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})$ satisfies a linear random partial differential equation (PDE) for $0\leq t\leq T$:
$\partial u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})=\underbrace{\mathcal{L}^{*}u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})\mathrm{d}t}_{\displaystyle\mathclap{\text{Forecast}}}+\underbrace{\mathcal{M}^{*}u_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x})\cdot\mathrm{d}\mathbf{x}}_{\displaystyle\mathclap{\text{Analysis}}},$ (1.8)
@@
$\mathcal{M}^{*}u_{t}(\mathbf{y}):=(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}\big(-\nabla_{\mathbf{y}}\cdot\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})u_{t}\big)+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1/2}\mathbf{f}^{\mathbf{x}}u_{t}\big),$ (1.10)
@@
$\overleftarrow{\partial u_{t}^{\text{{s}}}}(\mathbf{y}|\mathbf{x})=\underbrace{\mathcal{L}u_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})\mathrm{d}t}_{\displaystyle\mathclap{\text{Forecast}}}+\underbrace{\mathcal{M}u_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})\cdot\overleftarrow{\mathrm{d}\mathbf{x}}}_{\displaystyle\mathclap{\text{Analysis}}},$ (1.11)
@@
The diffusion interactions $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})$ are defined in the same manner as $\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}$, while $\overleftarrow{\partial u_{t}^{\text{{s}}}}$ and $\overleftarrow{\mathrm{d}\mathbf{x}}$ denote backward stochastic Itô integrals (i.e., the negative of the usual differentials up to the principal (linear) part).
@@ A.4.2 Managing the Influence of $\mathbf{x}_{\text{B}}$ on $\mathbf{y}(t)$’s Estimation
In Algorithm 2 we provide a brief outline of how to implement conditional ACI in pseudocode.
Data: Observations of $\mathbf{x}$ in $[0,T]$ (continuous or discrete), $p_{0}^{\text{f}}(\mathbf{y}|\mathbf{x})$
Result: $\mathcal{P}\big(p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})\big)$ for $t\in[0,T]$, used to assess $\big(\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$ at each time
1
2for $0\leq t\leq T$ do
    Use (1.1) and the observations of $\mathbf{x}=(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})^{\mathtt{T}}$ to compute $p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})$ by letting $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ during the corresponding analysis step of Bayesian data assimilation, defined by the forward/filter Kalman update operator $\mathcal{M}^{*}$ in (1.10);
    /*/ E.g., see Theorem B.1 and the proof of Theorem C.2 for the system in (2.1) as a concrete example
3
4 end for
5
6for $T\geq t\geq 0$ (moving backward) do
    Use $p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})$ and the observations of $\mathbf{x}=(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})^{\mathtt{T}}$ to compute $p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})$ by letting $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ during the corresponding analysis step of Bayesian data assimilation, defined by the backward/smoother Kalman update operator $\mathcal{M}$ in (1.12b);
    /*/ E.g., see Theorem B.2 and the proof of Theorem C.2 for the system in (2.1) as a concrete example
7
8 end for
9
10for $0\leq t\leq T$ do
    Calculate $\mathcal{P}\big(p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})\big)=\int p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})\log\left(\displaystyle\frac{p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})}{p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})}\right){\mathrm{d}}\mathbf{y}(t)$;
    /*/ Can also use any other type of information metric, e.g., a different $f$-divergence
11    if $\mathcal{P}\big(p^{\text{{s}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),p^{\text{{f}}|\mathbf{x}_{\text{{B}}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{{A}}})\big)>0$ then
12        $\big(\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$;
13
14   else
15       $\big(\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$;
16    end if
17
18 end for
Algorithm 2 Conditional ACI for the identification of $\big(\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$ over $t\in[0,T]$
@@ A.4.3 Arguments in Favor of the Conditional ACI Framework
The generalized ACI framework presented in Section A.4 and defined by (1.13)–(1.14) offers several key advantages for establishing (1.15):
@@
    Complete elimination of $\mathbf{x}_{\text{B}}$’s influence during data assimilation, ensuring any uncertainty reduction in $\mathbf{y}(t)$ stems solely from $\mathbf{x}_{\text{A}}$, consistent with our ACI objectives. Furthermore, during analysis, the correlations between $\mathbf{x}_{\text{A}}$ and $\mathbf{x}_{\text{B}}$ in their likelihood are necessarily sent to zero as the uncertainty of $\mathbf{x}_{\text{B}}$ goes to infinity. This further resolves the potential for spurious associations arising over the observable state space that can flow over when inferring causality from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$; see Theorem C.2. A rough analogue to this approach would be the assignment of an improper prior as an uninformative or diffuse prior in parametric Bayesian inference.
@@ A.4.4 Further Justification: Why Infinite Uncertainty Outperforms Marginalization
One might attempt to infer a conditional causal relationship from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$ by marginalizing over $\mathbf{x}_{\text{B}}(t)$:
$p_{t}\big(\mathbf{y}\big|\mathbf{x}_{\text{A}}(s\leq T)\big)=\int p_{t}\big(\mathbf{y},\mathbf{x}_{\text{B}}\big|\mathbf{x}_{\text{A}}(s\leq T)\big){\mathrm{d}}\mathbf{x}_{\text{B}}(t).$
@@
$\mathbf{y}$$\mathbf{x}_{\text{{B}}}$$\mathbf{x}_{\text{{A}}}$
For this causal chain, this marginalization (or conditioning) method might falsely suggest a direct causal link from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$. While these filter- and smoother-based distributions of $\mathbf{y}(t)$ given $\mathbf{x}_{\text{A}}$ may differ after marginalization (or conditioning), we cannot attribute causation to $\mathbf{x}_{\text{A}}$ under (1.4). The uncertainty reduction in $\mathbf{y}(t)$ inherently incorporates information transfer from $\mathbf{x}_{\text{B}}(t)$ due to their coupled dynamics, unlike the distributions in (1.13) where $\mathbf{x}_{\text{B}}$ trajectories are known but carry infinite uncertainty in their likelihood.
A.5 Dynamic Discovery of the Causal Influence Range (CIR)
The relative entropy between the smoother and filter solutions in (1.4) quantifies both the existence and strength of an assimilative causal link from $\mathbf{y}(t)$ to $\mathbf{x}$. Similarly, the generalized metric in (1.14) measures the conditional causal link $\big(\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$ and its intensity. However, these ACI metrics alone cannot determine the temporal extent of causal influence, namely, how many future values of $\mathbf{x}$ (or $\mathbf{x}_{\text{A}}$ conditioned on $\mathbf{x}_{\text{B}}$) are affected by $\mathbf{y}(t)$ over $[t,T]$. We define this temporal characteristic as the causal influence range (CIR), representing the future time window where $\mathbf{y}(t)$’s causal impact persists. Below we develop the theory for unconditional CIR; the conditional case (CCIR) follows analogously by substituting the appropriate equations with those from Section A.4.
In chaotic turbulent systems, new observations of $\mathbf{x}$ influence $\mathbf{y}(t)$’s estimation only within finite time windows. More chaotic dynamics accelerate memory decay, typically exhibiting exponential decay modulated by factors like $\mathbf{x}$’s signal-to-noise ratio [andreou2024adaptive]. Our framework estimates this CIR duration for $\mathbf{y}(t)\rightarrow\mathbf{x}$ by determining how many future $\mathbf{x}$ values are meaningfully affected.